### v1

@@ 1 Introduction
We investigate the impact of data quality issues from two code-based datasets, Devign and BigVul zhou2019devign ; fan2020ac . These datasets contain C/C++ functions from multiple projects and have known quality issues croft2023data . Our analysis reveals that these issues cause gradient instability during training and significantly impair a model’s capacity to capture complex patterns. Furthermore, our manual analysis of attention weights shows that data quality issues in the code-based data can reduce the code comprehension abilities of the deep learning model.
@@
We investigate the impact of data quality issues from two metric-based datasets, Openstack and QT, which frequently suffer from class imbalance problems mcintosh2018fix ; hoang2019deepjit . Our analysis reveals that quality issues in metric-based data lead to vanishing gradients and poor optimization during the training process. Additionally, we perform qualitative analysis using GradCAM, an explainable AI technique for visual analysis of model inputs. Our qualitative analysis demonstrates that models trained on imbalanced data focus on irrelevant tokens and struggle to generalize to unseen data.
@@
We evaluate the generalizability of our findings using six new datasets: D2A and Juliet (code-based), Spark and Mozilla (text-based), and Go and JDT (metric-based). Our analysis reveals that the data quality and processing issues in the training data lead to reduced model capacity, gradient instability, overfitting, and biased learning of the models, which align with our above findings. In contrast, the models trained on cleaned datasets show none of these issues. Such observations increase confidence in our findings and underscore the challenges of data bugs in deep learning models used in software engineering tasks.
@@ 3.2.1 Task Selection
Selecting representative tasks for each type of data above is a critical step in our analysis (Step 2, Fig. 1). By carefully choosing the tasks that use our selected data types, we can examine how data bugs manifest and affect the training of deep learning models. In particular, we have selected the following representative tasks for our analysis, as suggested by Yang et al. yang2022dlinse :
(a). Vulnerability Detection: Detection of vulnerabilities in software code is one of the key applications where deep learning models show promising results. These models often use source code or binary for their detection task.
(b). Duplicate Bug Report Detection: Duplicate bug report detection is a prominent use case of text-based data from software engineering that is leveraged by DL models. The task’s objective is to identify bug reports that describe the same underlying issue from previous bug reports.
(c). Defect Prediction: Predicting defects in code components using software metrics is a common application of deep learning in software engineering. This task involves predicting the defect proneness of software modules based on various code metrics (e.g., complexity, coupling, cohesion).
@@ 3.2.2 Baseline Selection
After selecting the representative tasks, we choose multiple baseline approaches for the selected tasks. These approaches provide baseline deep learning models, which are used in our subsequent analysis. To select baselines for each of the types of data, we used the following criteria: (a) the technique should use a neural network model to perform one of the three tasks above, and (b) the technique should provide a comprehensive replication package with source code, datasets, and instructions for reproducibility. Furthermore, we prioritized the baselines with clean and buggy versions in their replication packages. Based on these filtration criteria, we selected the following state-of-the-art baselines for our study.
(a) Code-based data: LineVul fu2022linevul for vulnerability detection. LineVul is a state-of-the-art transformer-based model for line-level vulnerability detection in code. The model takes code tokens as input and learns to predict the vulnerability status of each line.
(b) Text-based data: DCCNN he2020duplicate for duplicate bug report detection. DCCNN, short for Duplicate bug report detection using Convolutional Neural Networks, employs a CNN-based architecture to identify duplicate bug reports from their textual descriptions. The CNN model learns to extract relevant features and patterns from a pair of bug reports and then predict whether they are duplicates or not.
(c) Metric-based data: DeepJIT hoang2019deepjit for defect prediction. DeepJIT is an end-to-end deep learning framework for Just-In-Time (JIT) defect prediction. This technique computes software metrics from datasets of code changes and extracts salient features from commit messages. The learned features from commit messages and code changes are encoded into numerical matrices and then processed by separate CNN layers to predict whether the commit will likely introduce a defect.
Besides the relevance to the selected data types, these baselines have demonstrated state-of-the-art performance in their respective tasks, which justifies their selection for our study (Step 3, Fig. 1).
@@ 3.2.3 Datasets
Text-Based Eclipse lazar2014generating 74,376 bug reports Bug reports from Eclipse project (2001-2007)
Hadoop hadoop 14,016 bug reports Bug reports from Hadoop project (2006-2013)
@@ 3.3.2 Baseline Preparation
3. Baseline with Missing Preprocessing: For our third variant, we focused on preprocessing faults, which are among the most prevalent types of faults in deep learning, according to the existing taxonomy taxonomyRealFaults . We removed the preprocessing operations from the training pipeline in preparation for training (Step 6, Fig. 1). By setting up the model to simulate the missing preprocessing operation, we aim to understand how the absence of these crucial steps might impact a model’s performance and robustness.
@@ 3.3.3 Integrating Weights and Biases
To better understand the training process and model behaviour, we integrated Weights and Biases (W&B) logging wandb into the existing baselines and their variants (Step 7, Fig. 1). After successfully integrating W&B, we trained the models and monitored and captured their higher-level training, such as the model’s loss and accuracy during the training. Additionally, we logged lower-level statistics such as gradients, weights, and biases from all model layers. This allowed us to track a model’s progress and performance over time. Furthermore, we logged the runtime configurations to capture the computational environment in which the experiments were conducted. Finally, to account for the variability and stochasticity inherent in neural networks, we ran each baseline model and corresponding variants fifteen times arcuri2014hitchhiker .
@@ 3.5 Post-Hoc Analysis
To gain deeper insights into the impact of data bugs on our deep learning models, we conducted a post-hoc analysis using explainable AI techniques (Step 8, Fig. 1). These techniques help us understand how the models make predictions and identify the influential parts of the input data.
@@
For the code-based data, we analyzed LineVul’s attention weights. Since the LineVul model is attention-based, an analysis of its attention weights helps us identify the parts of the input code that the model focuses on when making predictions about vulnerabilities. By comparing the attention weights from models trained on clean and buggy data, we observed how data bugs affect a model’s attention and potentially lead to incorrect predictions.
@@
For the text-based data, we utilized t-SNE (t-Distributed Stochastic Neighbor Embedding) visualization van2008visualizing . t-SNE is a dimensionality reduction technique that helps us visualize high-dimensional data in a lower-dimensional space. By applying t-SNE to the representations learned by the DCCNN model at different layers, we visualized the decision boundary for duplicate and non-duplicate bug reports. This helps us understand how the learned representations differ between models trained on clean and buggy data, providing insights into the impact of data bugs on the model’s learning process.
@@
For the metric-based data, we implemented GradCAM selvaraju2017grad . In the context of defect prediction using DeepJIT, GradCAM helps us visualize which metrics and parts of the input features are most influential in the model’s decision. By contrasting the GradCAM outputs of models trained on clean data against those trained on buggy data, we identify how data bugs affect the model’s focus and potentially lead to incorrect defect predictions.

### v2

@@ 1 Introduction
We investigate the impact of data quality issues from two code-based datasets, Devign and BigVul zhou2019devign ; fan2020ac . These datasets contain C/C++ functions from multiple projects and have known quality issues croft2023data . Our analysis reveals that these issues cause gradient instability during training and significantly affect the weight and bias distribution of the deep learning models. Furthermore, our manual analysis of attention weights shows that data quality issues in the code-based data can reduce the code comprehension abilities of the deep learning model.
@@
We investigate the impact of data quality issues from two metric-based datasets, Openstack and QT, which frequently suffer from class imbalance problems mcintosh2018fix ; hoang2019deepjit . Our analysis reveals that quality issues in metric-based data lead to vanishing gradients and higher loss during the training process. Additionally, we perform qualitative analysis using GradCAM, an explainable AI technique for visual analysis of model inputs. Our qualitative analysis demonstrates that models trained on imbalanced data focus on irrelevant tokens and struggle to generalize to unseen data.
@@
We evaluate the generalizability of our findings using six new datasets: D2A and Juliet (code-based), Spark and Mozilla (text-based), and Go and JDT (metric-based). Our analysis reveals that the data quality and processing issues in the training data lead to near-zero biases, smaller weights, gradient instability, and overfitting, which align with our above findings. In contrast, the models trained on cleaned datasets show none of these issues. Such observations increase confidence in our findings and underscore the challenges of data bugs in deep learning models used in software engineering tasks.
@@ 2 Background
2.1 Data Bug
A data bug is a systematic error or flaw in the data pipeline that leads to incorrect, inconsistent, or misleading data, thereby compromising the integrity or performance of a machine learning or data-driven system islam2019fse . Formally, for a data pipeline $\mathcal{P}$ that transforms raw data $X_{\text{raw}}$ into processed data $X_{\text{proc}}$, a data bug exists if:
[…]
where $X_{\text{expected}}$ is the correct output of the pipeline under ideal conditions. Data bugs can arise from issues such as data collection from unreliable sources, faulty transformations, incorrect feature engineering, erroneous dimensionality reduction, and inappropriate data type conversions that alter the fundamental nature of the data. They may also arise from missing preprocessing steps, such as failure to handle missing values, overlooking duplicate entries, and neglecting the inconsistencies in the data.
Class Imbalance: Class imbalance refers to a scenario where the distribution of classes in the dataset is highly skewed. Formally, given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ with $y_{i}\in\{1,2,...,K\}$, class imbalance exists if the class prior probabilities $P(Y=k)$ satisfy:
$\exists\ k,l\in\{1,...,K\}\ \text{such that}\ \frac{P(Y=k)}{P(Y=l)}\geq\tau,$
where $\tau\gg 1$ is a threshold indicating significant disparity (e.g., $\tau=10$ for severe imbalance). For instance, in software engineering domains, moderate imbalance is characterized by $\tau\approx 1.5$ (corresponding to ratios like 60:40) and might be acceptable. However, severe imbalance occurs at $\tau\geq 3$ (corresponding to ratios of 75:25 or greater), which is undesirable mahbub2023defectors ; li2022robust .
Concept Drift: Concept drift refers to a change in the underlying data distribution over time, specifically in the relationship between input features $X$ and the target variable $Y$. For a time-varying distribution $P_{t}(X,Y)$, concept drift occurs between times $t_{1}$ and $t_{2}$ if:
$P_{t_{1}}(Y\mid X)\neq P_{t_{2}}(Y\mid X),$
even if $P_{t_{1}}(X)=P_{t_{2}}(X)$. This distinguishes it from covariate shift (where only $P(X)$ changes).
Label Noise: Label noise arises when training labels are corrupted. For a true labeling function $f:\mathcal{X}\rightarrow\mathcal{Y}$, observed labels $\tilde{y}_{i}$ are perturbed such that:
$\tilde{y}_{i}=\begin{cases}y_{i}&\text{with probability }1-\eta,\\ z\sim\mathcal{Y}\setminus\{y_{i}\}&\text{with probability }\eta,\end{cases}$
where $\eta\in[0,1]$ is the noise rate frenay2014comprehensive . Label noise can be uniform (random flips) or structured (biased toward specific classes).
Missing Preprocessing Operations: Missing preprocessing operations occur when input data deviates from the expected distribution assumed by a deep learning (DL) model due to omitted transformations. Let $T:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ denote a required preprocessing function (e.g., normalization, imputation). If raw data $X$ is fed to a model trained on $T(X)$, the input distribution mismatch degrades the model’s performance. Formally, the model $M$ optimized for $P(T(X))$ faces a distribution shift when applied to $P(X)$, violating the i.i.d. assumption. In other words, the missing preprocessing operations introduce a shift in the distribution between the data the model was trained on and the data it encounters during inference. This shift in the distribution violates the assumption that the distributions during training and inference will be identical, and this causes a drop in the model’s performance.
@@ 2.2 Data Quality Issues in Software Engineering Datasets
2.3 Preprocessing Faults in Deep Learning Systems
In this section, we discuss the most common types of preprocessing faults in deep learning systems, as per the existing taxonomy by Humbatova et al. taxonomyRealFaults . These faults broadly manifest in two categories: missing preprocessing and incorrect preprocessing.
Missing Preprocessing: Missing preprocessing refers to a scenario when a preprocessing step leading to better performance has not been applied at all. As suggested by the existing literature taxonomyRealFaults , missing preprocessing can significantly impact the model performance. For code analysis, cases were reported where the lexer and parser failed to catch certain code structures due to missing preprocessing steps. In text processing, particularly with social media data and software engineering text, issues arose from not handling special characters and different encodings (UTF-8, Unicode). With metric-based data, the absence of correlation analysis between features led to redundant features in the training data, and missing normalization for numerical values affected model performance.
Incorrect Preprocessing: Incorrect preprocessing refers to a scenario when a preprocessing step was either inappropriate or incorrect. According to existing literature taxonomyRealFaults , incorrect preprocessing can manifest in subtle ways. In code preprocessing, incorrect token replacement and lexer implementations led to improper parsing of programming constructs. For text data, developers encountered problems with incorrect handling of special tokens, where preprocessing tokens were incompatible with the existing toolchain. One specific case involved the Stanford Parser producing incorrect tags for certain words, leading to the wrong interpretation of the data. For metric-based features, when different features of the numerical data were scaled inconsistently across different ranges (e.g., [0,1] versus [-1,1]), this resulted in performance degradation taxonomyRealFaults .
In this study, we focus exclusively on missing preprocessing faults for several reasons. First, missing preprocessing faults occur more frequently in deep learning systems than incorrect preprocessing taxonomyRealFaults . Second, missing preprocessing can be systematically studied by removing the preprocessing steps. However, there currently exists no systematic approach for simulating wrong preprocessing, as existing mutation techniques focus primarily on training, model, and data faults humbatova2021deepcrime . Third, manually selecting the incorrect preprocessing operations could introduce bias, leading to unreliable findings.
@@ 3.2.1 Task Selection
Selecting representative tasks for each type of data above is a critical step for our analysis (Step 2, Fig. 1). By carefully choosing the tasks that use our selected data types, we can examine how data bugs manifest and affect the training of deep learning models. In particular, we have selected classification tasks for our analysis due to their several benefits. First, classification-based tasks tackle many important challenges in software engineering (e.g., vulnerability detection, bug localization, duplicate bug report detection, code clone detection), leveraging different types of data. On the other hand, generation tasks in software engineering are primarily focused on code yang2022dlinse . Thus, selecting generation tasks could limit our ability to study the impact of data quality and preprocessing issues across different types of data. Second, classification tasks provide straightforward evaluation metrics (e.g., accuracy, precision, recall) for measuring the impact of data quality issues, which is crucial for our analysis. Thus, we have selected the following representative tasks for our analysis, as suggested by Yang et al. yang2022dlinse :
(a) Vulnerability Detection: Detection of vulnerabilities in software code is one of the key applications where deep learning models show promising results. These models often use source code or binary for their detection task.
(b) Duplicate Bug Report Detection: Duplicate bug report detection is a prominent use case of DL models leveraging text-based data from software engineering. The task’s objective is to identify the incoming bug reports that describe the same underlying issue from the past bug reports.
(c) Defect Prediction: Predicting defects in code components (e.g., classes, methods) using software metrics is a common application of deep learning in software engineering. The task involves predicting the defect proneness of software modules based on various code metrics (e.g., complexity, coupling, cohesion).
@@ 3.2.2 Baseline Selection
After selecting the representative tasks, we choose multiple baseline approaches for them. These approaches provide baseline deep learning models, which are used in our subsequent analysis. To select baselines for each task, we used the following criteria: (a) the technique should use a neural network model to perform one of the three tasks above, and (b) the technique should provide a comprehensive replication package with source code, datasets, and instructions for reproducibility. We also look for baselines with clean and buggy versions of the data in their replication packages. Based on these filtration criteria, we selected the following state-of-the-art baselines for our study.
(a) Code-based data: LineVul fu2022linevul , CodeBERT ding2024vulnerability : LineVul and CodeBERT are state-of-the-art transformer-based models for line-level vulnerability detection in code. The models take code tokens as input and learn to predict the vulnerability status of each line.
(b) Text-based data: DCCNN he2020duplicate , BERT-MLP messaoud2022dbrd : DCCNN, short for Duplicate bug report detection using Convolutional Neural Networks, employs a CNN-based architecture to identify duplicate bug reports leveraging their textual descriptions. The CNN model first learns to extract relevant features and patterns from a pair of bug reports and then predicts whether they are duplicates. BERT-MLP, a representation learning-based technique, uses the BERT model to learn the contextual relationship between words. Then, it uses a softmax-based classifier to determine if two bug reports are duplicates or not based on their representation.
(c) Metric-based data: DeepJIT hoang2019deepjit , CodeBERTJIT guo2023jit : DeepJIT, and CodeBERTJIT are end-to-end deep learning frameworks for Just-In-Time (JIT) defect prediction. These techniques compute software metrics from datasets of code changes and extract salient features from commit messages. DeepJIT uses CNN to extract the salient features, whereas CodeBERTJIT utilizes CodeBERT to extract the features. The learned features from commit messages and code changes are encoded into numerical matrices and then processed by separate CNN layers to predict whether the commit will likely introduce a defect.
Besides strong relevance to our selected data types, these baseline techniques have demonstrated state-of-the-art performance in their respective tasks, which justifies their selection for our study (Step 3, Fig. 1).
@@ 3.2.3 Datasets
Text-Based Eclipse lazar2014generating 74,376 bug reports Bug reports from Eclipse project (2001-2014)
Hadoop hadoop 14,016 bug reports Bug reports from Hadoop project (2012-2014)
@@
3.2.4 Quality Issues in Existing Datasets
The benchmark datasets utilized in our analysis exhibit several quality concerns that warrant discussion:
  • Vulnerability Detection (Devign, BigVul): Prior research has identified significant quality issues in these datasets croft2023data . Existing analysis indicates that only 54% and 80% of the samples in Big-Vul and Devign maintain correct labels, demonstrating substantial label noise croft2023data . Furthermore, these datasets face challenges with temporal relevance, as 15% of Devign samples and 25% of Big-Vul samples no longer reflect the current trends of vulnerability croft2022data .
  • Duplicate Bug Report Detection (Hadoop, Eclipse): These datasets exhibit similar quality concerns regarding data representation and temporal validity. Studies have shown that only 5.2% of Eclipse bug reports and 2.7% of Hadoop bug reports are marked as duplicates zhang2023duplicate . Besides, the majority of their bug reports were submitted in 2014 or earlier, potentially leading to concept drift problems in these datasets zhang2023duplicate . Research has also demonstrated that approximately 25-40% of bug reports in Bugzilla-based systems contain incorrect labels, indicating substantial label noise in these systems tantithamthavorn2015impact .
  • Defect Prediction (OpenStack and QT): The OpenStack and QT datasets present multiple quality challenges. Research has revealed that their dataset construction methodology may introduce significant labelling errors, potentially affecting up to 55% of the data yatish2019icse . These datasets also exhibit major class imbalance, with defective lines constituting only 20% of the total samples hoang2019deepjit . Additionally, these datasets were constructed using commits before 2014, which might introduce significant concept drift in the techniques leveraging these datasets hoang2019deepjit .
@@ 3.3.2 Baseline Preparation
3. Baseline with Missing Preprocessing: For our third variant, we focused on preprocessing faults, which are among the most prevalent types of faults in deep learning, according to the existing taxonomy taxonomyRealFaults . We removed the preprocessing operations from our training pipeline dealing with different types of data (Step 6, Fig. 1): for code-based data, we omitted the stop word removal, line separation, and sequence length normalization; for text-based data, we excluded stemming, stop word removal, and case conversion; and for metric-based data, we eliminated data normalization and feature scaling. By setting up the model to simulate the missing preprocessing operation, we aim to understand how the absence of these crucial steps might impact a model’s performance and robustness.
@@ 3.3.3 Integrating Weights and Biases
To better understand the training process and model behaviour, we integrated Weights and Biases (W&B) logging wandb into the existing baselines and their variants (Step 7, Fig. 1). W&B’s built-in logging capabilities and predefined hooks facilitated the extraction of gradients, weights, and biases. After configuring W&B to watch the model parameters, it automatically tracked these metrics and attached hooks to each layer, eliminating the need for manual implementation. During training, W&B captured the full distribution of gradients flowing through each layer, weight matrices, and bias vectors. The hooks tracked gradient changes during backpropagation, providing insights into the learning dynamics of different layers. These metrics were stored in W&B’s dashboard for detailed analysis of the model’s internal dynamics throughout the training process. We implemented this integration across all model variants to maintain consistency in our experimental methodology. Additionally, we logged runtime configurations to document the computational environment. To account for the variability and stochasticity inherent in neural networks, we ran each baseline model and corresponding variants fifteen timesarcuri2014hitchhiker .
@@ 3.4 Quantitative Analysis
3.5 Determining the Symptoms
We developed a data-driven approach to identify symptoms of data quality and preprocessing issues in deep learning models. Our methodology consists of four key steps:
First, we established baseline parameter distributions by training 15 models on clean, properly preprocessed data. For each model layer, we collected descriptive statistics of weights, biases, and gradients during training. In particular, we capture the mean, median, maximum, minimum, variance, standard deviation, skewness, kurtosis and sparsity of these measures. These statistics characterize the expected behaviors of properly trained models and serve as our reference point.
Second, we trained multiple models on datasets containing known preprocessing and data quality issues. Similar to the first step, we captured descriptive statistics of weights, biases and gradients.
Third, we aggregated these statistics layer-by-layer across multiple models to identify common patterns. We then compared the aggregated statistics between the bug-free and buggy models to detect deviations in the buggy model’s statistics, using the bug-free model as our reference. We also flagged outliers in the buggy model’s statistics using three complementary statistical methods:
  • Z-Score analysis for weights and biases: Computing mean ($\mu$) and standard deviation ($\sigma$) identify data points with high Z-score, $|Z|>3$ (i.e., $|X-\mu|>3\sigma$tukey1977exploratory .
  • Interquartile Range (IQR) for gradients: Calculating 25th ($Q_{1}$) and 75th ($Q_{3}$) percentiles to identify values outside $[Q_{1}-1.5\times IQR,Q_{3}+1.5\times IQR]$ leys2013detecting .
  • Distribution shape analysis: Measuring skewness ($>1$) and kurtosis
    ($|kurtosis|>3$) to detect distribution distortions westfall2014kurtosis .
By capturing the above statistics from model parameters, we attempted to detect consistent statistical patterns that could serve as reliable indicators of data bugs in the DL models.
Finally, we analyzed statistics and averaged them across models and layers where these patterns appeared consistently in the buggy models. We then compared these against our bug-free models, which provided essential reference distributions of normal parameter behavior. These bug-free models were crucial for establishing reliable statistical thresholds and understanding what constitutes expected parameter distributions in properly trained models. By contrasting the buggy models’ statistics against these reference distributions, we could precisely identify which deviations represented genuine data quality issues rather than acceptable variations. These thresholds were also validated against validation datasets to confirm their effectiveness in detecting known data quality issues. This process resulted in a set of reliable statistical criteria that can help identify symptoms and their prevalence in model training.
3.6 Determining the Impact on Downstream Performance
After determining the symptoms of data quality and preprocessing issues, we determine the impact of these symptoms on the performance of the downstream task. To determine how individual symptoms affect model performance while controlling for confounding factors; we performed a systematic analysis of models exhibiting isolated symptoms. For each symptom identified through our quantitative analysis, we filtered our experimental runs to obtain a subset of models that displayed only that particular symptom, excluding any models that exhibited multiple concurrent symptoms. This filtering process ensured that our analysis focused solely on the direct relationship between individual symptoms and performance degradation. We then aggregated performance metrics (accuracy, precision, recall, F1-score) across all models that uniquely exhibited each type of symptom. By analyzing models with isolated symptoms, we could attribute performance changes to specific symptoms without the confounding effects of symptom co-occurrence.
@@ 3.7 Post-Hoc Analysis
To gain deeper insights into the impacts of data bugs on our deep learning models, we conducted a post-hoc analysis using Explainable AI techniques (Step 8, Fig. 1). These techniques help us identify the influential parts of the input data and understand how the models make predictions.
@@
For the code-based data, we analyzed LineVul’s attention weights. Since the LineVul model is attention-based, an analysis of its attention weights helps us identify the parts of the input code that the model focuses on when making predictions about vulnerabilities. By comparing the attention weights from models trained on clean and buggy data, we observed how data bugs affect a model’s attention and potentially lead to incorrect predictions. Given an attention weight tensor $A^{l}\in\mathbb{R}^{h\times n\times n}$ from layer $l$, which contains attention weights for $h$ heads and $n$ tokens, we first average the attention weights across all heads. For each token pair $(j,k)$ in the attention matrix, we compute the mean attention weight $\bar{A}^{l}_{jk}$ using:
$\bar{A}^{l}_{jk}=\frac{1}{h}\sum_{i=1}^{h}A^{l}_{i,j,k}$ (1)
This gives us a single $n\times n$ attention matrix (for n tokens) that represents the average attention weights across all heads. To quantify the overall importance of each token $j$ in the sequence, we sum up all the attention weights directed towards that token using:
$score_{j}=\sum_{i=1}^{n}\bar{A}^{l}_{ij}$ (2)
This produces a scalar $score_{j}$ for each token $j$, representing how much attention that token receives from all other tokens on average. When calculated for all tokens, this results in a vector of importance scores $[score_{1},...,score_{n}]$, which measures each token’s relative importance in the context.
@@
For the text-based data, we utilized t-SNE (t-Distributed Stochastic Neighbor Embedding) visualization van2008visualizing . t-SNE is a dimensionality reduction technique that helps us visualize high-dimensional data in a lower-dimensional space. By applying t-SNE to the representations learned by the DCCNN model he2020duplicate at different layers, we visualized the decision boundary for duplicate and non-duplicate bug reports. Given high-dimensional embeddings $X={x_{1},...,x_{N}}$ and their low-dimensional counterparts $Y={y_{1},...,y_{N}}$, t-SNE minimizes their distribution gap by employing the Kullback-Leibler divergence:
$KL(P||Q)=\sum_{i}\sum_{j}p_{ij}\log\frac{p_{ij}}{q_{ij}}$ (3)
where $p_{ij}$ and $q_{ij}$ are the pairwise similarities of the duplicate bug reports in the high and low-dimensional spaces, respectively. In the high-dimensional space, similarities are computed using a Gaussian distribution:
$p_{ij}=\frac{\exp(-||x_{i}-x_{j}||^{2}/2\sigma^{2})}{\sum_{k\neq i}\sum_{l\neq k}\exp(-||x_{k}-x_{l}||^{2}/2\sigma^{2})}$ (4)
In the low-dimensional space, similarities are computed using a Student’s t-distribution with one degree of freedom:
$q_{ij}=\frac{(1+||y_{i}-y_{j}||^{2})^{-1}}{\sum_{k\neq i}\sum_{l\neq k}(1+||y_{k}-y_{l}||^{2})^{-1}}$ (5)
The t-SNE based visualization above helps us understand how the learned representations differ between models trained on clean and buggy data, providing insights into the impact of data bugs on the model’s learning process.
@@
In the case of metric-based data, we implemented GradCAM selvaraju2017grad for our analysis. In the context of defect prediction using DeepJIT, GradCAM helps us visualize which metrics and parts of the input features are most influential in the model’s decision, as per the existing literature issta2021deepjit . For a given class $c$, let $y^{c}$ be the score for class $c$ before softmax. The GradCAM weights $\alpha^{c}_{k}$ for each feature map $k$ are computed as:
$\alpha^{c}_{k}=\frac{1}{Z}\sum_{i}\sum_{j}\frac{\partial y^{c}}{\partial A^{k}_{ij}}$ (6)
where $A^{k}$ represents the $k$-th feature map in the network, which corresponds to learned representations of different aspects of our input metrics. Unlike image applications where $(i,j)$ refers to spatial positions, in our metric-based data, $i$ indexes the samples (code changes) and $j$ indexes the feature dimensions (software metrics). The term $Z$ normalizes the gradients across all input instances.
The final GradCAM heatmap is obtained through a weighted combination of feature maps followed by a ReLU activation:
$L^{c}_{GradCAM}=ReLU\left(\sum_{k}\alpha^{c}_{k}A^{k}\right)$ (7)
This equation performs a linear combination of the feature maps $A^{k}$, where each map is weighted by its importance coefficient $\alpha^{c}_{k}$. The ReLU function (Rectified Linear Unit) then sets all negative values to zero, ensuring that only features positively contributing to the class prediction are highlighted in the visualization. The resulting heatmap $L^{c}_{GradCAM}$ has the same dimensions as our feature matrix and indicates which metrics and feature combinations most strongly influence the model’s classification decision for class $c$.
By contrasting the GradCAM outputs of models trained on clean data against those trained on buggy data, we identify how data bugs affect the model’s attention to different metrics and potentially lead to incorrect defect predictions.